OpenStudy (anonymous):

Ψ(x,t) =Ax/a, if 0<=x<=b , =A(b-x)/(b-a), if a <=x<=b, =0, otherwise, whr a,b are constant...1.sketch Ψ(x,0) as a fuction of x

OpenStudy (michele_laino):

Hint: I think that your function is defined as below: \[\Large \psi \left( x \right) = \left\{ {\begin{array}{*{20}{c}} {A\frac{x}{a},\quad 0 \leqslant x \leqslant a} \\ \begin{gathered} \hfill \\ A\frac{{b - x}}{{b - a}},\quad a \leqslant x \leqslant b \hfill \\ \hfill \\ 0,\quad otherwise \hfill \\ \end{gathered} \end{array}} \right.\] In such way, the wave function is continuous at point \(x=a\) Now, before to draw the graph of such function, we have to normalize it, with this condition: \[\Large \int_{ - \infty }^{ + \infty } {{{\left| {\psi \left( x \right)} \right|}^2}} dx = 1\] after a simple computation, we get: \[\Large A = \sqrt {\frac{3}{b}} \] therefore, we have to draw the graph of this function: \[\Large \psi \left( x \right) = \left\{ {\begin{array}{*{20}{c}} {\sqrt {\frac{3}{b}} \left( {\frac{x}{a}} \right),\quad 0 \leqslant x \leqslant a} \\ \begin{gathered} \hfill \\ \sqrt {\frac{3}{b}} \left( {\frac{{b - x}}{{b - a}}} \right),\quad a \leqslant x \leqslant b \hfill \\ \hfill \\ 0,\quad otherwise \hfill \\ \end{gathered} \end{array}} \right.\]